# Benchmarking Ultra-Low-Power NPUs

### v2

@@ Abstract.
Efficient on-device neural network (NN) inference has various advantages over cloud-based processing, including predictable latency, enhanced privacy, greater reliability, and reduced operating costs for vendors. This has sparked the recent rapid development of microcontroller-scale NN accelerators, often referred to as neural processing units (NPUs), designed specifically for ultra-low-power applications.
In this paper, we present the first comparative evaluation of a number of commercially-available NPUs, as well as the first independent benchmarks for several of these platforms. We develop and open-source a model compilation framework to enable consistent benchmarking of quantized models across diverse NPU hardware. Our benchmark targets end-to-end performance and includes model inference latency, power consumption, and memory overhead, alongside other factors. The resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including NPUs exhibiting unexpected scaling behaviors with increasing model complexity. Our framework provides a foundation for further evaluation of NPU platforms alongside valuable insights for both hardware designers and software developers in this rapidly evolving space.
@@ 1. INTRODUCTION
Performing neural network (NN) inference on constrained devices has applications across numerous domains, including wearable health monitoring (wearablehealth, ), smart agriculture (agriculture, ), real-time audio processing (audio, ), and predictive maintenance (predmaintenance, ). On-device inference offers various advantages over cloud-based alternatives: improved latency for time-critical applications, enhanced privacy, as well as reduced operating costs for vendors, by eliminating the need to transmit sensitive data, and improved reliability by removing dependence on network connectivity. Given their unique form factor and low power consumption, microcontrollers (MCUs) are widely used in resource-constrained environments. However, their performance is often constrained by limitations in memory capacity, throughput, and compute.
@@
performance data centers to ultra-low-power and embedded devices. At the resource-constrained end of the spectrum, microcontroller-scale neural processing units (NPUs) have recently emerged, designed to operate within extremely tight power envelopes — in the milliwatt or sub-milliwatt range — while still providing low latency for real-time inference. These devices represent a new class of accelerator, combining the power efficiency of MCUs with the cognitive capabilities previously exclusive to more powerful computing platforms. The core advantage of NPUs stems from their ability to exploit the inherent parallelism of neural networks with dedicated multiply-accumulate (MAC) arrays alongside specialized memory structures for weight storage. Such architectural specialization enables NPUs to achieve orders of magnitude improvement in latency compared to general-purpose MCUs executing equivalent workloads.
Despite the growing availability of NPU platforms, the field lacks a standardized evaluation or comprehensive benchmark suite. Existing benchmarks focus solely on Analog Devices’ MAX78000 (max78000, ; moss2022ultra, ; clay2022benchmarking, ), lacking any side-by-side comparison with other platforms. Hardware vendors provide performance metrics based on proprietary evaluation frameworks, often using disparate NN models, quantization strategies, and other varying optimizations. This heterogeneity across evaluation methods, and absence of independent verification of vendor-provided performance claims, creates uncertainty for hardware designers and embedded software developers in selecting the most suitable NPU platform for their application’s constraints. The lack of standardized benchmarking also hampers research by obscuring the relationship between architectural design and real-world performance. Given the rapid pace of development and increasing diversity of available NPU platforms, establishing reliable comparative benchmarks has become an urgent need for the field.
@@
Independent Benchmarks: We also provide the first fine-grained and independent performance benchmarks for several NPU platforms that have not previously been subject to third-party evaluation, offering unbiased verification of vendor performance claims.
Open-Source Model Compilation Framework: We develop and release111Upon publication an open-source framework that enables consistent and simplified transplanting of NN models across diverse NPU platforms, reducing the engineering overhead associated with cross-platform evaluation.
Developer Recommendations: Informed by our benchmark results, we provide actionable recommendations to developers regarding platform selection, key focus areas for model optimization, and trade-offs for various application scenarios and constraints.
In developing a unified compilation and benchmarking framework, we standardize model representations across the various NPU platforms, enabling direct comparison of latency, memory, and energy performance. Our evaluation also includes fine-grained analysis of the various model execution stages, from NPU initialization and memory input/output overheads to CPU pre/post-processing – aspects that can significantly impact end-to-end performance but are often not addressed in technical evaluations. The resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including NPUs exhibiting unexpected scaling behaviors with increasing model complexity. We hope our findings provide valuable insights to both developers and hardware architects.
@@ 2. BACKGROUND & MOTIVATION
2.1. Resource-Constrained
Neural Computing
@@ 2.1. Resource-Constrained
MCUs are compact, low-power computing platforms, often reliant on a single CPU and shared memory bus (saha2022machine, ). While MCUs are commonly adopted for resource-constrained IoT applications (lin2022device, ; kwon2023tinytrain, ; huang2024low, ), they generally lack the computational resources for efficient NN inference. Specifically, the computational capability of typical MCUs is often limited to a few million MAC operations per second, far below the tens of billions MACs/s required for real-time NN inference. Their absence of dedicated hardware acceleration results in large latency overheads and elevated power consumption during NN processing. Limited SRAM and flash memory also often poses challenges for efficiently managing the large weight matrices required for NN models.
@@
microcontroller-scale NPUs are emerging as a response. These specialized NN accelerators offer dedicated neural processing hardware, providing higher throughput for NN workloads, meeting the stringent requirements of real-time NN inference (manor2022custom, ; wang2022bed, ; song2024tada, ) while maintaining low-power operation. Collectively, NPUs position themselves as a key solution for efficient, real-time NN processing in resource-constrained environments.
@@ 2.2. NPU Hardware Design
NPU hardware design is optimized for efficient tensor operations via specialized MAC units and parallelizable memory hierarchies (caronti2023fine, ; gong2025dex, ). Fig. 1 illustrates the architecture of a typical NPU, composed of a systolic array of processing elements (PEs). Notably, each PE contains its own MAC units and, importantly, its own weight memory space to avoid memory contention and maximize parallelization. The array of PEs is linked by an inter-PE communication grid, which connects to a large global buffer and SRAM/DRAM via an on-chip network (sze, ). Efficient memory hierarchy optimization is achieved by partitioning available RAM, along with implementing high-bandwidth memory interfaces and data prefetching mechanisms, addressing the memory bottlenecks faced by traditional MCUs when handling large NN model weights. NPUs mainly vary by their number of PEs, PE layout and clustering, memory hierarchy layout, and the availability/amount of storage/MAC units in each PE.
@@ 2.3. Benchmarking NPU Platforms
Adoption of NPU Platforms: The increasing demand for on-device neural computing has accelerated the development and commercialization of NPUs. This is evidenced by the growing number of vendors, including Arm (arm_ethos_u, ), who have released NPU platforms to market.
@@
Limitations of Existing Benchmarks: Existing benchmarks of NPUs focus on a single platform, lacking horizontal comparisons across the now wide variety of available platforms (moss2022ultra, ; clay2022benchmarking, ; huang2024energy, ). This narrow perspective limits understanding of the variations in performance and task-based applicability across different NPUs. Existing standalone benchmarks also have significant shortcomings. Chiefly, most focus solely on the model’s inference forward pass, overlooking other adjacent operations within the end-to-end model inference or application pipeline(s), such as NPU initialization, memory input/output (I/O), and CPU pre/post-processing. While often neglected, these factors can significantly impact overall performance and efficiency.
@@ 3. INFRASTRUCTURE & METHODOLOGY
We begin by detailing our benchmark hardware and models, then provide a comprehensive overview of our benchmarking framework and model inference pipeline.
@@ 3.1. Hardware
To provide a comprehensive benchmark, we evaluate a diverse range of widely-used, commercially-available NPU platforms, from ultra-low-power NPUs to high-performance NPU-equipped system-on-chip (SoC) architectures. These are evaluated alongside MCUs without dedicated neural hardware for comparison. Our selection covers a wide range of computational capabilities (¡5 to ¿500 GOPs), memory configurations (128 KB to 2 MB RAM), and bit-width support (1-bit quantized to 32-bit floating-point operations). Fig. 2 provides a visualization of peak GOPs (Giga Operations Per Second) vs. peak power for the various NPU platforms included in our benchmark (on a log scale). Table 1 details our set of benchmark NPUs, and we provide more detail on each platform below.
@@
We also include the MILK-V Duo (milkv, ), a RISC-V SoC built around the CVITEK CV1800B processor. Unlike the previous MCUs/NPUs, it runs a Linux OS, supporting more flexible NN workloads at a much-increased power budget. This platform represents the upper bound of our evaluation in terms of computational power and software flexibility.
| MCU | CPU(s) | NPU | Flash | RAM | GOPs | Bit Cap. |
@@
| NXP-MCXN947 | Cortex-M33 x2 | eIQ Neutron | 2 MB | 512 KB | 4.8 | 8 |
@@
| MILK-V | RISC-V XuanTie C906 x2 | CV1800B | - | 64 MB | 500 | 8, 16, 32 |
@@ 3.1.1. Note on CPU Frequency
We configure the various NPU platforms to operate at a uniform CPU frequency. While this permits direct comparison of architectural efficiency, it should be noted that many of the platforms are capable of operating at higher frequencies than evaluated – approaching the GHz range in some cases. Our method intentionally isolates architectural efficiency, but further experimentation could explore the impact of varying CPU frequencies on end-to-end latency and power consumption.
Other hardware parameters are largely standardized by default. Further work could investigate their impact on performance; the number of active PEs, memory layout configuration, and other hardware-specific optimizations are variably configurable across the NPU platforms and can influence overall efficiency.
@@ 3.2. Models
CIFAR10-NAS: the optimal CNN model generated by neural architecture search (NAS) for the CIFAR-10 dataset, combining diverse convolutional units, pooling layers, and unique connectivity patterns. The model was generated using the Once-for-All (OFA) NAS framework, a weight-sharing-based framework that decouples search and training by constructing a supernet model from which various hardware-specific subnet models can be derived (onceforall, ). This is our largest model, with 74.3 Million MACs (MMACs) and 36.4 Million FLOPs (MFLOPs). Trained on the CIFAR-10 dataset, with 3x32x32 input size and 10-class output.
ResidualNet: a CNN framework built around residual functions, helping to mitigate gradient vanishing. ResidualNet has 37.7 MMACs and 18.5 MFLOPs. Trained on the CIFAR-100 dataset, with 3x32x32 input size and 100-class output.
SimpleNet: a simpler CNN framework composed of a basic stack of convolutional and pooling layers, without complex branches or residual functions. Despite its basic architecture, SimpleNet often outperforms more complex models including certain versions of ResidualNet (simplenet, ). SimpleNet has 38.0 MMACs and 18.5 MFLOPs. Trained on the CIFAR-100 dataset, with 3x32x32 input size and 100-class output.
Autoencoder: a symmetric encoder-decoder model. This model is our simplest, with just 0.5 MMACs and 0.2 MFLOPs. Trained on a machine fault detection dataset, generated using the SpectraQuest Machinery Fault Simulator (MotorDatasets, ). The input/output size is 3x256.
YoloV1: a single-stage object detection CNN that uses multi-scale output feature maps to predict bounding boxes and class probabilities. YoloV1 has 43.83 MMACs and 21.2 MFLOPs. Trained for person-only detection on the COCO dataset (coco, ), with input size 3x96x96 and 3 output layers, which result from pruning its final layers for cross-platform uniformity (more details below). The non-max suppression (NMS) post-processing step is performed on CPU.
Note: MMACs/MFLOPs are forward-only.
| Model | Input Shape | Output Shape | MMACs | MFLOPs | |
| CIFAR10-NAS | 3x32x32 | 1x10 | 74.2512 | 36.3776 | |
| ResidualNet | 3x32x32 | 1x100 | 37.7812 | 18.4612 | |
| SimpleNet | 3x32x32 | 1x100 | 38.0006 | 18.4612 | |
| Autoencoder | 3x256 | 3x256 | 0.5455 | 0.2020 | |
| YoloV1 | 3x96x96 | 1x12x12x12 1x12x12x2 1x12x12x10 | 43.8294 | 21.2244 |
@@ 3.2.3. Compilation
Fig 3 below details our model compilation framework for converting a base (Torch/ONNX/TFLM) model into various platform-specific formats. Our framework will be released as open-source, and we hope its use can ease the process of cross-platform model compilation and benchmarking.

### v3

@@ Abstract.
Efficient on-device neural network (NN) inference offers predictable latency, improved privacy and reliability, and lower operating costs for vendors than cloud-based inference. This has sparked recent development of microcontroller-scale NN accelerators, also known as neural processing units (NPUs), designed specifically for ultra-low-power applications.
We present the first comparative evaluation of a number of commercially-available NPUs, including the first independent benchmarks for multiple platforms. To ensure fairness, we develop and open-source a model compilation pipeline supporting consistent benchmarking of quantized models across diverse microcontroller hardware. Our resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including certain NPUs exhibiting unexpected scaling behaviors with model complexity. This work provides a foundation for ongoing evaluation of NPU platforms, alongside offering practical insights for both hardware and software developers in this rapidly evolving space.
@@ 1. INTRODUCTION
Performing neural network (NN) inference on constrained devices has applications across numerous domains, including wearable health monitoring (wearablehealth, ), smart agriculture (agriculture, ), real-time earable audio processing (audio, ), and predictive maintenance (predmaintenance, ). Embedding inference also offers improved privacy over cloud-based alternatives, by eliminating the need to transmit sensitive data, alongside improved latency for time-critical applications, reduced operating costs for vendors, and improved reliability by removing dependence on network connectivity. Given their unique form factor and low power consumption, microcontrollers (MCUs) are widely used in resource-constrained environments. However, their performance is also often constrained by limitations in memory capacity, throughput, and compute.
@@
performance data centers to ultra-low-power and embedded devices. At the resource-constrained end of the spectrum, microcontroller-scale neural processing units (NPUs) have recently emerged, designed to operate within extremely tight power envelopes – in the milliwatt or sub-milliwatt range – while still providing latency sufficient to support real-time inference. These platforms represent a new class of accelerator, combining the power efficiency of MCUs with the cognitive capabilities previously exclusive to more powerful computing platforms. The core advantage of NPUs stems from their ability to exploit the inherent parallelism of neural networks with dedicated multiply-accumulate (MAC) arrays alongside specialized memory structures for weight storage. This architectural specialization enables NPUs to achieve orders of magnitude improvement in latency compared to general-purpose MCUs executing equivalent workloads.
Despite the growing availability of NPU platforms, the field lacks a standardized and comprehensive evaluation or benchmark suite. Existing benchmarks focus solely on Analog Devices’ MAX78000 (max78000, ; moss2022ultra, ; clay2022benchmarking, ), lacking side-by-side comparison with other platforms. Naturally, hardware vendors themselves provide performance metrics, but these are usually based on proprietary evaluation frameworks, using disparate NN models, quantization strategies, and other varying optimizations. This heterogeneity across evaluation methods, and absence of independent verification of vendor-provided performance claims, creates uncertainty for hardware designers and embedded software developers in selecting the most suitable platform for their application’s constraints. The lack of standardized benchmarks also hampers research by obscuring the relationship between architectural design and inference performance. Given the rapid pace of development and increasing diversity of available NPU platforms, establishing reliable comparative benchmarks has become an urgent need.
@@
Independent Platform Benchmarks: We also provide the first fine-grained and independent performance benchmarks for several NPU platforms that have not previously been subject to third-party evaluation, offering unbiased verification of vendor performance claims.
Open-Source Model Compilation Toolchain: We develop and release111https://github.com/j0shmillar/uNPU-Bench an open-source toolchain to support consistent and simplified transplanting of NN models across our NPU platforms, reducing the engineering overhead associated with cross-platform evaluation.
Recommendations for Developers: Informed by our benchmark results, we provide actionable recommendations to developers regarding platform selection, key focus areas for model optimization, and trade-offs for various applications and constraints.
In developing a unified compilation and benchmarking pipeline, we standardize model representations across the various NPU platforms, enabling direct comparison of latency, memory, and energy performance. The evaluation also includes fine-grained analysis of the various stages of model execution, from NPU initialization and memory input/output overheads to CPU pre/post-processing – aspects that can significantly impact end-to-end performance but are often not addressed in technical evaluations. The resulting analysis uncovers both expected performance trends as well as surprising disparities between hardware specifications and actual performance, including certain NPUs exhibiting unexpected scaling behaviors with increasing model complexity. We hope our findings provide valuable insights to both developers and hardware architects.
@@ 2. BACKGROUND & MOTIVATION
2.1. Constrained Neural Computing
@@
MCUs are compact, low-power computing platforms, often reliant on a single CPU and shared memory bus (saha2022machine, ). While MCUs are commonly adopted for resource-constrained IoT applications (lin2022device, ; kwon2023tinytrain, ; huang2024low, ), they generally lack the computational resources for efficient NN inference. Specifically, the computational capability of typical MCUs is often limited to a few million MAC operations per second, far below the tens of billions MACs/s required for real-time NN inference. Their absence of dedicated hardware acceleration results in large latency overheads and elevated power consumption during NN processing. Limited SRAM and flash memory also often poses challenges for efficiently managing the large weight matrices required for NN inference.
@@
microcontroller-scale NPUs are emerging as a response. These specialized NN accelerators offer dedicated neural processing hardware, providing higher throughput for NN workloads, meeting the stringent requirements of real-time NN inference (manor2022custom, ; wang2022bed, ; song2024tada, ) while maintaining low-power operation. Collectively, NPUs position themselves as a key solution for real-time NN processing in low-power environments.
@@ 2.2. NPU Hardware Design
NPU hardware design is optimized for efficient tensor operations via specialized MAC units and parallelizable memory hierarchies (caronti2023fine, ; gong2025dex, ). Fig. 1 illustrates the architecture of a typical NPU, composed of a systolic array of processing elements (PEs). Notably, each PE contains its own MAC units and, importantly, its own weight memory space to avoid memory contention and maximize parallelization. The array of PEs is linked by an inter-PE communication grid, which connects to a large global buffer and SRAM/DRAM via an on-chip network (sze, ). Efficient memory hierarchy optimization is achieved by partitioning available RAM, along with implementing high-bandwidth memory interfaces and data prefetching mechanisms, addressing the memory bottlenecks faced by traditional MCUs when handling large NN weights. NPUs mainly vary by their number of PEs, PE layout and clustering, memory hierarchy layout, and the availability/amount of storage/MAC units in each PE.
@@ 2.3. Benchmarking NPU Platforms
Limitations of Existing Benchmarks: Existing benchmarks of NPUs focus on a single platform, lacking horizontal comparisons across the now wide variety of available platforms (moss2022ultra, ; clay2022benchmarking, ; huang2024energy, ). This narrow perspective limits understanding of the variations in performance and applicability across different NPUs. Existing standalone benchmarks also have significant shortcomings. Chiefly, most focus solely on the model’s inference forward pass, overlooking other adjacent operations within the end-to-end model inference or application pipeline(s), such as NPU initialization, memory input/output (I/O), and CPU pre/post-processing. While often neglected, these factors can significantly impact overall performance and efficiency.
@@ 3.1. Hardware
To provide a comprehensive benchmark, we evaluate a diverse range of widely-used, commercially-available NPU platforms, from ultra-low-power NPUs to NPU-equipped system-on-chip (SoC) architectures. These are evaluated alongside MCUs without dedicated neural hardware for comparison. Our selection covers a wide range of computational capabilities (¡5 to ¿500 GOPs), memory configurations (128 KB to 2 MB RAM), and bit-width support (1-bit quantized to 32-bit floating-point operations). Fig. 2 provides a visualization of peak GOPs (Giga Operations Per Second) vs. peak power for the various NPU platforms included in our benchmark (on a log scale). Table 1 details our set of benchmark NPUs, and we provide more detail on each platform below.
@@
We also include the MILK-V Duo (milkv, ), a RISC-V SoC built around the CVITEK CV1800B processor. Unlike the previous MCUs/NPUs, it can run Linux variants or an RTOS, such as FreeRTOS, supporting more flexible NN workloads at a much-increased power budget. This platform represents the upper bound of our evaluation in terms of computational power and software flexibility.
| MCU | CPU(s) | NPU | Flash | RAM | GOPs (max) | Bit Cap. |
@@
| NXP-MCXN947 | Cortex-M33 (2) | eIQ Neutron | 2 MB | 512 KB | 4.8 | 8 |
@@
| MILK-V | RISC-V XuanTie C906 (2) | CV1800B | - | 64 MB | 500 | 8, 16, 32 |
@@ 3.1.1. Note on CPU Frequency
We configure the various NPU platforms to operate at a uniform CPU frequency. While this permits direct comparison of architectural efficiency, it should be noted that many of the platforms are capable of operating at higher frequencies than evaluated – approaching the GHz range in some cases. Our method intentionally isolates architectural efficiency, but further experimentation could explore the impact of varying CPU frequencies on end-to-end latency and power consumption. All other hardware parameters are fixed to platform-default settings to ensure comparability. These include the number of active processing elements (PEs) or compute cores, on-chip SRAM banking configuration, and any vendor-specific accelerator tuning parameters (, tiling sizes, DMA burst length, cache enablement). Memory layout settings – such as channel ordering – are likewise determined by the compiler defaults for each platform. While these defaults yield a “fair” cross-platform baseline, they may not represent peak achievable performance.
@@ 3.2. Models
CIFAR10-NAS: the optimal memory-constrained CNN model for the CIFAR-10 dataset, generated using the Once-for-All (OFA) NAS framework, a weight-sharing-based framework that decouples search and training by constructing a supernet model from which various hardware-specific subnet models can be derived (onceforall, ). . It combines diverse convolutional units, frequent 1×1 convolutions for channel mixing, alternating pooling layers, and irregular channel scaling – patterns that are characteristic of hardware-aware NAS designs. This is our largest model, with 74.3 Million MACs (MMACs) and 36.4 Million FLOPs (MFLOPs). Trained on the CIFAR-10 dataset, with 3x32x32 input size and 10-class output.
SimpleNet: a simpler CNN framework composed of a basic stack of convolutional and pooling layers (simplenet, ). SimpleNet has 38.0 MMACs and 18.5 MFLOPs. Trained on the CIFAR-100 dataset, with 3x32x32 input size and 100-class output.
ResidualNet: a SimpleNet variant built around residual functions, helping to mitigate gradient vanishing and introducing a non-trivial scheduling problem for inference compilers. ResidualNet and SimpleNet are structurally matched in size and compute, but differ in connectivity, enabling direct measurement of skip-connection overheads. ResidualNet has 37.7 MMACs and 18.5 MFLOPs. Trained on the CIFAR-100 dataset, with 3x32x32 input size and 100-class output.
AI8XAutoEnc: encoder–decoder architecture adapted from the AI8X framework (i.e., for MAX78000/2). It combines two initial 1D convolutional layers with a series of fully connected layers, including a strong bottleneck compression down to 4 latent features before reconstruction. The decoder uses small FC expansions rather than mirroring the encoder’s convolutional structure. This model is our simplest, with just 0.5 MMACs and 0.2 MFLOPs. Trained on a machine fault detection dataset, generated using the SpectraQuest Machinery Fault Simulator (MotorDatasets, ). The input/output size is 3x256.
YoloV1: a compact, single-stage object detection CNN, with its final layers pruned for uniformity across platforms. This network is deep (20 convolutional layers) but narrow (maximum 32 channels), with heavy use of 1×1 reductions and rapid spatial downsampling via max pooling. YoloV1 has 43.83 MMACs and 21.2 MFLOPs. Trained for person-only detection on the COCO dataset (coco, ), with input size 3x96x96. The network produces multi-scale output feature maps for bounding boxes and class probabilities, requiring CPU-side non-max suppression (NMS).
| Model | Input | Output | #Params | #Layers | Conv Depth | Max Conv Ch. | FC Depth | Peak Act. RAM (MB, B=1)† | Lifetime Pressure (MBsteps)‡ | MMACs | MFLOPs | |||
| CIFAR10-NAS | 33232 | 110 | 298,762 | 11 | 10 | 128 | 1 | 0.07 | 18.39 | 74.25 | 36.38 | |||
| ResidualNet | 33232 | 1100 | 383,012 | 14 | 14 | 512 | 0 | 0.14 | 2.98 | 37.78 | 18.46 | |||
| SimpleNet | 33232 | 1100 | 383,012 | 14 | 14 | 512 | 0 | 0.14 | 2.45 | 38.00 | 18.46 | |||
| AI8XAutoEnc | 3256 | 3256 | 136,800 | 7 | 2 | 128 | 5 | 0.07 | 2.89 | 0.55 | 0.20 | |||
| YoloV1 | 39696 |
40,700 | 20 | 20 | 32 | 0 | 0.02 | 0.80 | 43.83 | 21.22 |
† Peak activation RAM: maximum live activation footprint during the forward pass (batch=1, INT8), excluding parameter storage.
‡ Lifetime Pressure: across all tracked activations, where counts the number of subsequent leaf-module steps over which activation must remain live until its last use. Higher values indicate reduced opportunities for buffer reuse.
@@ 3.2.3. Compilation
Fig 3 below details our model compilation toolchain for converting a base (Torch/ONNX/TFLM) model into various platform-specific formats. We open-source our toolchain222https://github.com/j0shmillar/uNPU-Bench and hope its use can ease the process of cross-platform model compilation and benchmarking.